# What is the nature of Little Red Dots and what is not, MIRI SMILES edition

### v1

@@ Abstract
We study little red dots (LRD) detected by JADES and covered by the SMILES MIRI survey. Our sample contains 31 sources, % detected in the two bluest MIRI bands, 40% in redder filters. The median/quartiles redshifts are (55% spectroscopic). We analyze the rest-frame ultraviolet through near/mid-infrared spectral energy distributions of LRDs combining NIRCam and MIRI observations, using a variety of modeling techniques that include emission from stars, dust, and (un)obscured active galactic nuclei (AGN). The NIRCamMIRI colors, for m, are bluer than direct pure emission from AGN tori; the spectral slope flattens in the rest-frame near-infrared, consistent with a 1.6 m stellar bump. Both observations imply that stellar emission makes the dominant contribution at these wavelengths, expediting a stellar mass estimation: the median/quartiles are . The number density of LRDs is Mpc, accounting for % of the global population of galaxies with similar redshifts and masses. The flat ultraviolet spectral range is dominated by young stars. The rest-frame near/mid-infrared (2–4 m) spectral slope reveals significant amounts of dust (bolometric stellar attenuation mag) heated by strong radiation fields arising from highly embedded compact sources. Our models imply kpc heating knots, containing dust-enshrouded OB stars or an AGN producing a similar radiation field, obscured by mag. We conclude that LRDs are extremely intense and compact starburst galaxies with mass-weighted ages 5–10 Myr, very efficient in producing dust, their global energy output dominated by the direct and dust-recycled emission from OB stars, with some contribution from obscured AGN in the mid-infrared.
@@ 1 Introduction
In the very first month of JWST science operations, Labbé et al. (2023a) identified a sample of compact sources with distinct spectral energy distributions (SEDs) presenting two clear breaks (Lyman and Balmer) in the NIRCamHST data. The SEDs were also characterized by a change of slope: they had red colors at observed wavelengths between 2 and (at least) m, the range covered by the NIRCam long-wavelength (LW) channels, and a flat SED in the short wavelength (SW) bands. Labbé et al. (2023a) identified these SW-blueLW-red sources as very massive, M, maybe significantly-obscured, mag, galaxies at . They have been claimed to represent a new galaxy population revealed through the new capabilities provided by JWST, easily reaching magnitudes down to 28-29 mag from 1 to 5 m.
@@
This high-mass interpretation was challenged by adopting models with prominent nebular emission, which could account for their red nature due to the non-negligible contribution of emission lines (and nebular continuum) to the broad-band fluxes as they enter the NIRCam passbands for different redshifts, implying significantly smaller (by a factor of 10 or more) stellar masses (see Endsley et al., 2023a, b; Pérez-González et al., 2023a; Desprez et al., 2023). The estimated masses of early galaxies would also be reduced significantly if they had top-heavy stellar initial mass functions (Woodrum et al., 2023; Wang et al., 2023). A simpler explanation in some cases is redshift errors. Indeed, Kocevski et al. (2023) presented spectroscopy for one of the claimed massive galaxies in Labbé et al. (2023a), showing an overestimation of the photometric redshift which could partly explain the high mass value for this and some other sources, as opposed to the more accurate photometric redshift and (10 times) smaller mass found in Pérez-González et al. (2023a) for this same source. However, the general agreement among different authors in the determination of photometric redshifts for this type of source implies that redshift errors are not worrisome in a statistical sense (estimations agree well between papers for % of the samples).
@@
Jointly with their colors, the extreme compactness of these objects led to the term Little Red Dots (LRDs, as first suggested by Matthee et al. 2023b). Understanding LRDs has become more complex given that there is a variety of selection techniques that arrive at very similar types of objects in terms of compact morphology and red LW-to-SW colors, with high levels of overlap but also ”contamination” from other types of sources (e.g., not so compact red galaxies or no little not-so-red dots). Labbé et al. (2023a) originally looked for double-break sources, which resulted in a flat or blue SED at wavelengths shorter than m and a very red one at longer wavelengths to 5 m (SW-blueLW-red SEDs). Eventually arriving at similar objects, Barro et al. (2023), Akins et al. (2023), or Labbé et al. (2023b) used colors such as F277W-F444W in their selection, adding additional constraints in the SW channels. Some LRDs also entered in the selection carried out by Pérez-González et al. (2023a), Barrufet et al. (2023) and Williams et al. (2023a) based on F150W-F356W and F150W-F444W (see also red sources selected with F200W-F444W in Rodighiero et al. 2023 and Bisigello et al. 2023), enclosed on larger investigations about the nature and cosmic relevance of “HST-dark” sources. Finally, complementary to photometric techniques, spectroscopic data have also been able to identify LRDs, mainly based on the selection of strong H or [OIII] emitters in blind NIRCam/grism spectroscopic surveys (Matthee et al., 2023a, b).
Beyond selection, concerning their nature, some LRDs (at the 10–20% level) have been confirmed to be brown dwarfs based on NIRCam colors, proper motions, and spectroscopy (Hainline et al., 2023a; Langeroodi & Hjorth, 2023; Burgasser et al., 2023).
@@
The question about the general nature of LRDs was further tangled by finding evidence for the presence of Active Galactic Nuclei (AGN). Kocevski et al. (2023) identified a broad (1000 km s) component in the H emission of one galaxy in the Labbé et al. (2023a) sample, which implied the presence of a 10 M super-massive black hole (SMBH). This large SMBH mass is consistent with the discoveries of high-redshift AGN in other studies, all indicating that black holes grew rapidly in the early Universe (Díaz-Santos et al., 2021; Larson et al., 2023; Maiolino et al., 2023a; Matthee et al., 2023b; Übler et al., 2023; Harikane et al., 2023; Natarajan et al., 2023; Bogdán et al., 2023), many lying above the Magorrian et al. (1998) relationship (Pacucci et al., 2023; Stone et al., 2023). Further spectroscopic analyses identified AGN in some additional LRDs and stellar emission in others, even both at different wavelengths in the same LRD (Matthee et al., 2023b; Greene et al., 2023; Killi et al., 2023). ALMA non-detections have also been used to defend the AGN nature of the bulk of the LRD emission (Labbé et al., 2023b).
@@
Solving the challenges outlined above requires going longer into the mid-infrared were stellar emission and AGN can be distinguished. MIRI data at observed wavelengths shorter than 8 m for a handful of LRDs was presented in Akins et al. (2023), Barro et al. (2023, also with a F1000W detection), indicating that they could be either dusty starbursts (see also Rodighiero et al. 2023) or obscured AGN, no definitive proof could be presented with the bluest MIRI bands. Williams et al. (2023a) used longer wavelength MIRI data from the SMILES program and found that the averaged SEDs (stacked in observed bands) of the LRDs flattened beyond 5 m, interpreting this behavior as the expected turnover of a normal stellar SED at rest m. In addition, and at odds with the AGN interpretation of the ALMA data of LRDs in the A2744 field presented Labbé et al. 2023b, Williams et al. (2023a) found that even deeper ALMA observations of LRDs in GOODS-S would agree with SED models mostly dominated by stars (with some implications for the global dust emission SED shape).
@@
In this paper, we investigate the nature of LRDs using the best multi-wavelength, large area MIRI data available to date, those coming from the SMILES survey (Lyu et al., 2023). MIRI, in fact, is essential to disentangle the nature of LRDs. If dominated by star formation, the SEDs of LRDs should show a 1.6 m bump, redshifted to m at (e.g. Williams et al., 2023a). On the other hand, an obscured AGN will yield a steep SED. The significant uncertainty in the relative importance of star formation and nuclear activity needs both spectroscopic (e.g., Kocevski et al., 2023; Killi et al., 2023) and photometric data to cover the widest spectral range possible, as well as a careful SED analysis of a statistically representative sample. In modeling these sources, we will also explore the uncertainties by using four different sets of software with a variety of approaches and assumptions.
@@
Throughout the paper, we assume a flat cosmology with , and a Hubble constant . We use AB magnitudes (Oke & Gunn, 1983). All stellar mass and SFR estimations assume a universal Chabrier (2003) IMF, which is a very relevant assumption for LRDs (and might not be adequate), given the results presented in this paper (about stellar masses and inferred stellar ages).
@@ 2.1 NIRCam-based selection
The sample of LRDs analyzed in this paper has been gathered from the JWST Advanced Deep Extragalactic Survey, JADES (Eisenstein et al., 2023a) Data Release 2 catalogs (Eisenstein et al. 2023b), which also gathered data from FRESCO (Oesch et al., 2023) and JEMS (Williams et al., 2023b). We also used the Rieke et al. (2023) catalog for spectroscopic redshifts, which included values measured with data taken by NIRCam (Oesch et al., 2023) and NIRSpec (Bunker et al., 2023). Given the small (nearly point-like) nature of our sources of interest and the fact that eventually we wanted to analyze SEDs including MIRI data, we used 0.25″ radius PSF-matched photometry as our fiducial aperture for selection, also checking results using smaller radii, more adequate for NIRCam. We searched for NIRCam SW-blueLW-red sources defined by F277W-F444W1 mag and F150W-F200W0.5 mag colors, and magnitudes F444W28 mag, following the strategies presented in Barro et al. (2023), Labbé et al. (2023b) and Greene et al. (2023). The two latter references further use a concentration criterion to select compact sources, but we found no need for it after applying the F150W-F200W0.5 mag cut. In any case, the typical concentration of our original sample is F(F444W)/F(F444W) (median and quartiles). Given that the ratio is close to 1 and it has been obtained after applying an aperture correction, we conclude that most of the sources are nearly point-like.
@@
We further restrict the sample to the area covered by MIRI, arriving at a sample of 37 galaxies. As shown by Hainline et al. (2023a), this type of color selection of high redshift galaxies might be contaminated by brown dwarfs. To account for that, and following their analysis, we removed from the original sample the 4 sources with color F115W-F150W–0.5 mag. Two other LRDs were removed since they were identified with brown dwarfs with proper motion (Hainline et al., 2023a). Our final sample contains 31 galaxies detected in the SMILES 34.9 arcmin area, i.e., the surface density is 0.9 LRD arcmin. Postage stamps in RGB format constructed with NIRCam data are shown in Figure 2. Table 1 provides all relevant information about the selection of our sample of LRDs.
@@ 2.2 Mid-infrared properties
In this section, we discuss the detections of our sample of LRDs in the MIRI data gathered by the Systematic Mid-infrared Instrument Legacy Extragalactic Survey, SMILES (Lyu et al., 2023). We first describe briefly this dataset, and then discuss the detection fractions in MIRI as a function of wavelength.
@@ 2.2.1 MIRI data description
The MIRI data gathered by the SMILES survey (program ID 1207, PI: G. Rieke, Lyu et al. 2023) were reduced with jwst pipeline version 1.12.3, reference files in pmap version 1138, which includes improved absolute photometric calibration and aperture corrections based on a better characterization of the PSFs, especially in the shortest wavelengths.
@@
Photometry in the MIRI bands was performed in circular apertures with radii 0.2″, 0.3″, 0.4″, and 0.5″, applying the corresponding aperture corrections to each measurement. Uncertainties were calculated following the procedure explained in Pérez-González et al. (2023b), which gathers non-contiguous (i.e., independent) pixels to avoid the effect of correlated noise. Different measurements for our sample of LRDs agreed within 0.1 mag and we eventually chose for each band the weighted mean of all fluxes, typically the one corresponding to 70% encircled energy for each band. The photometry was revised visually to avoid contamination from cosmic ray showers, which we found to affect one source. We only kept fluxes with S/N5 and considered 5 upper limits otherwise.
@@ 2.2.2 MIRI detections
Out of the 31 LRDs selected with the JADES NIRCam data down to F444W28 mag, we detect [19,22,12,7,4,2,1,0] galaxies, translating to detection fractions (in %) [61,71,39,23,13,7,3,0] at levels [26.1,25.8,24.7,24.3,24.2,23.0,22.6,20.8] mag (measured in circular aperture of radius equal to the PSF FWHM, corrected to infinite extent apertures) for filters [F560W,F770W,F1000W,F1280W,F1500W,F1800W, F2100W,F2550W].
The detection fractions depend strongly on F444W magnitude. For the brightest galaxies, F444W26.5 mag, the detection fractions are (in %) [93,100,73,47,20,13,7,0] in [F560W,F770W,F1000W,F1280W,F1500W,F1800W, F2100W,F2550W], respectively.
We divided the sample of LRDs into 4 subsets detected up to a given MIRI band: (1) 5 galaxies detected in F1280W and longer wavelengths, hereafter the Golden Five; (2) 7 galaxies detected only up to F1000W (i.e., excluding the previous type); (3) 7 sources detected only up to F770W (excluding the previous types); (4) and, lastly, the 12 galaxies not detected in any MIRI band. The panels and histograms in Figure 1, as well as the postage stamps in Figure 2, show these subsets with different colors. The first figure illustrates that the MIRI detection fraction depends strongly on the F444W and it is nearly independent of the F150W-F200W color. Interestingly, the F277W-F444W color does exhibit some differences between subsets, with the Golden Five being among the bluest, and the F1000W sample among the reddest (we will comment on this in following sections). It is worth mentioning that only two of the Golden Five galaxies would be identified as LRDs under a more restrictive color selection of F277W-F444W1.5 mag.

### v2

@@ Abstract
We study 31 little red dots (LRD) detected by JADES/NIRCam and covered by the SMILES/MIRI survey, of which % are detected in the two bluest MIRI bands and 40% in redder MIRI filters. The median/quartiles redshifts are (55% spectroscopic). The spectral slopes flatten in the rest-frame near-infrared, consistent with a 1.6 m stellar bump but bluer than direct pure emission from active galactic nuclei (AGN) tori. The apparent dominance of stellar emission at these wavelengths for many LRDs expedites stellar mass estimation: the median/quartiles are . The number density of LRDs is Mpc-3, accounting for % of the global population of galaxies with similar redshifts and masses. The rest-frame near/mid-infrared (2–4 m) spectral slope reveals significant amounts of warm dust (bolometric attenuation mag). Our spectral energy distribution modeling implies the presence of kpc diameter knots, heated by either dust-enshrouded OB stars or an AGN producing a similar radiation field, obscured by mag. We find a wide variety in the nature of LRDs. However, the best-fitting models for many of them correspond to extremely intense and compact starburst galaxies with mass-weighted ages 5–10 Myr, very efficient in producing dust, with their global energy output dominated by the direct (in the flat rest-frame ultraviolet and optical spectral range) and dust-recycled emission from OB stars with some contribution from an obscured AGN (in the infrared).
@@ 1 Introduction
In the very first month of JWST science operations, Labbé et al. (2023a) identified a sample of compact sources with distinct spectral energy distributions (SEDs) presenting two clear breaks (Lyman and Balmer) in the NIRCamHST data. The SEDs were also characterized by a change of slope: they had red colors at observed wavelengths between 2 and (at least) m, the range covered by the NIRCam long-wavelength (LW) channels, and a flat SED in the short wavelength (SW) bands. Labbé et al. (2023a) identified these SW-blueLW-red sources as very massive, M⊙, maybe significantly-obscured, mag, galaxies at . They have been claimed to represent a new galaxy population revealed through the new capabilities provided by JWST, easily reaching magnitudes down to 28-29 mag from 1 to 5 m.
@@
This high-mass interpretation was challenged by adopting models with prominent nebular emission, which could account for their red nature due to the non-negligible contribution of emission lines (and nebular continuum) to the broad-band fluxes as they enter the NIRCam passbands for different redshifts, implying significantly smaller (by a factor of 10 or more) stellar masses (see Endsley et al., 2023a, b; Pérez-González et al., 2023a; Desprez et al., 2023). The estimated masses of early galaxies would also be reduced significantly if they had top-heavy stellar initial mass functions (Woodrum et al., 2023; Wang et al., 2023).
A simpler explanation in some cases is redshift errors or mis-identifications. Indeed, Kocevski et al. (2023) presented spectroscopy for one of the claimed massive galaxies in Labbé et al. (2023a), showing an overestimation of the photometric redshift which could partly explain the high mass value for this and some other sources, as opposed to the more accurate photometric redshift and (10 times) smaller mass found in Pérez-González et al. (2023a) for this same source. However, the general agreement among different authors in the determination of photometric redshifts for this type of source implies that redshift errors are not worrisome in a statistical sense (estimations agree well between papers for % of the samples). In addition, some purported LRDs (at the 10–20% level) have been found to be brown dwarfs based on NIRCam colors, proper motions, and spectroscopy (Hainline et al., 2023a; Langeroodi & Hjorth, 2023; Burgasser et al., 2023).
@@
Jointly with their colors, the extreme compactness of the bona fide high-redshift objects led to the term Little Red Dots (LRDs, as first suggested by Matthee et al. 2023b). Understanding LRDs has become more complex given that there is a variety of selection techniques that arrive at very similar types of objects in terms of compact morphology and red LW-to-SW colors, with high levels of overlap but also “contamination” from other types of sources (e.g., not so compact red galaxies or little not-so-red dots). Labbé et al. (2023a) originally looked for sources with SED breaks, which resulted in a flat or blue SED at wavelengths shorter than m and a very red one at longer wavelengths to 5 m (SW-blueLW-red SEDs). Eventually arriving at similar objects, Barro et al. (2023), Akins et al. (2023), or Labbé et al. (2023b) used colors such as F277W-F444W in their selection, adding additional constraints in the SW channels. Some LRDs also entered in the selection carried out by Pérez-González et al. (2023a), Barrufet et al. (2023) and Williams et al. (2023a) based on F150W-F356W and F150W-F444W (see also red sources selected with F200W-F444W in Rodighiero et al. 2023 and Bisigello et al. 2023), which were in some cases included in larger investigations about the nature and cosmic relevance of “HST-dark” sources. Finally, complementary to purely photometric techniques, spectroscopic data with less comprehensive photometry have also been able to identify LRDs, mainly based on the selection of strong H or [OIII] emitters in blind NIRCam/grism spectroscopic surveys (Matthee et al., 2023a, b).
@@
The question about the general nature of LRDs was further tangled by finding evidence for the presence of Active Galactic Nuclei (AGN). Kocevski et al. (2023) identified a broad (1000 km s-1) component in the H emission of one galaxy in the Labbé et al. (2023a) sample, which implied the presence of a 107 M⊙ super-massive black hole (SMBH). This large SMBH mass is consistent with the discoveries of high-redshift AGN in other studies, all indicating that black holes grew rapidly in the early Universe (Díaz-Santos et al., 2021; Larson et al., 2023; Maiolino et al., 2023a; Matthee et al., 2023b; Übler et al., 2023; Harikane et al., 2023; Natarajan et al., 2023; Bogdán et al., 2023), many lying above the Magorrian et al. (1998) relationship (Pacucci et al., 2023; Stone et al., 2023). Further spectroscopic analyses identified AGN in some additional LRDs and stellar emission in others, even both at different wavelengths in the same LRD (Matthee et al., 2023b; Greene et al., 2023; Killi et al., 2023). ALMA non-detections have also been used to defend the AGN nature of the bulk of the LRD emission (Labbé et al., 2023b).
@@
Solving the challenges outlined above requires going further into the mid-infrared where stellar emission and AGN can be distinguished. MIRI data at observed wavelengths shorter than 8 m for a handful of LRDs was presented in Akins et al. (2023), Barro et al. (2023, also with a F1000W detection), indicating that they could be either dusty starbursts (see also Rodighiero et al. 2023) or obscured AGN, but no definitive conclusions could be reached with just the bluest MIRI bands. Williams et al. (2023a) used longer wavelength MIRI data from the SMILES program and found that the averaged SEDs (stacked in observed bands) of the LRDs flattened beyond 5 m, interpreting this behavior as the expected turnover of a normal stellar SED at rest m. In addition, and at odds with the AGN interpretation of the ALMA data of LRDs in the A2744 field presented by Labbé et al. 2023b, Williams et al. (2023a) found that even deeper ALMA observations of LRDs in GOODS-S could agree with SED models mostly dominated by stars (with some implications for the global dust emission SED shape).
@@
In this paper, we investigate the nature of LRDs using the best multi-wavelength, large area MIRI data available to date, those coming from the SMILES survey (Lyu et al., 2023). MIRI, in fact, is essential to disentangle the nature of LRDs. If dominated by star formation, the SEDs of LRDs should show a 1.6 m bump, redshifted to m at (e.g. Williams et al., 2023a). On the other hand, an obscured AGN will yield a steep SED. The significant uncertainty in the relative importance of star formation and nuclear activity needs both spectroscopic (e.g., Kocevski et al., 2023; Killi et al., 2023) and photometric data to cover the widest spectral range possible, as well as a careful SED analysis of a statistically representative sample. In modeling these sources, we will also explore the uncertainties by using four different sets of modeling software with a variety of approaches and assumptions.
@@
Throughout the paper, we assume a flat cosmology with , and a Hubble constant . We use AB magnitudes (Oke & Gunn, 1983). All stellar mass and SFR estimations assume a universal Chabrier (2003) IMF, which is a very relevant assumption for these parameters (but might not be accurate given the extreme conditions).
@@ 2.1 NIRCam-based selection
The sample of LRDs analyzed in this paper has been gathered from the JWST Advanced Deep Extragalactic Survey, JADES (Eisenstein et al., 2023a) Data Release 2 catalogs (Eisenstein et al. 2023b), which also gathered data from the First Reionization Epoch Spectroscopically Complete Observations (FRESCO, Oesch et al. 2023) and the JWST Extragalactic Medium-band Survey (JEMS, Williams et al. 2023b). We also used the Rieke et al. (2023) catalog for spectroscopic redshifts, which included values measured with data taken by NIRCam (Oesch et al., 2023) and NIRSpec (Bunker et al., 2023). Given the small (nearly point-like) nature of our sources of interest and the fact that eventually we wanted to analyze SEDs including MIRI data, we used 0.25″ radius PSF-matched photometry as our fiducial aperture for selection, but also checked the results using smaller radii, more suitable for NIRCam only. We searched for NIRCam SW-blueLW-red sources defined by F277W-F444W1 mag and F150W-F200W0.5 mag colors, and magnitudes F444W28 mag, following the strategies presented in Barro et al. (2023), Labbé et al. (2023b) and Greene et al. (2023). The two latter references further use a concentration criterion to select compact sources, but we found no need for it after applying the F150W-F200W0.5 mag cut. In any case, the typical concentration of our original sample is indicated by the flux ratio (after applying the appropriate aperture corrections) of F(F444W)r=0.25″/F(F444W)r=0.10″ (median and quartiles). That is, most of the sources are nearly point-like.
@@
We further restrict the sample to the area covered by MIRI within the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES, Lyu et al. 2023), arriving at a sample of 37 galaxies. As shown by Hainline et al. (2023a), this type of color selection of high redshift galaxies might be contaminated by brown dwarfs. To account for that, and following their analysis, we removed from the original sample the 4 sources with color F115W-F150W–0.5 mag. Two other LRDs were removed since they were identified with brown dwarfs with proper motion (Hainline et al., 2023a). Our final sample contains 31 galaxies detected in the SMILES 34.9 arcmin2 area, i.e., the surface density is 0.9 LRD arcmin-2. Postage stamps in RGB format constructed with NIRCam data are shown in Figure 2. Table 1 provides all relevant information about the selection of our sample of LRDs.
@@ 2.2 Mid-infrared properties
In this section, we discuss the detections of our sample of LRDs in the MIRI data gathered by SMILES (Lyu et al., 2023). We first describe briefly this dataset, and then discuss the detection fractions in MIRI as a function of wavelength.
@@ 2.2.1 MIRI data description
The MIRI data gathered by the SMILES survey (program ID 1207, PI: G. Rieke, Lyu et al. 2023) were reduced with JWST pipeline version 1.12.3, reference files in pmap version 1138, which includes improved absolute photometric calibration and aperture corrections based on a better characterization of the PSFs, especially at the shortest wavelengths.
@@
Photometry in the MIRI bands was performed in circular apertures with radii of 0.2″, 0.3″, 0.4″, and 0.5″, applying the corresponding aperture corrections to each measurement. Uncertainties were calculated following the procedure explained in Pérez-González et al. (2023b), which gathers non-contiguous (i.e., independent) pixels to avoid the effect of correlated noise. Different measurements for our sample of LRDs agreed within 0.1 mag and we eventually chose for each band the weighted mean of all fluxes, typically the one corresponding to 70% encircled energy for each band. The photometry was revised visually to avoid contamination from cosmic ray showers, which we found to affect one source. We only kept fluxes with S/N5 and considered 5 upper limits otherwise.
@@ 2.2.2 MIRI detections
Out of the 31 LRDs selected with the JADES NIRCam data down to F444W28 mag, we detect [19,22,12,7,4,2,1,0] galaxies for MIRI bands [F560W,F770W,F1000W,F1280W,F1500W,F1800W, F2100W,F2550W] respectively. The corresponding detection fractions (in %) are [61,71,39,23,13,7,3,0] at levels of [26.1,25.8,24.7,24.3,24.2,23.0,22.6,20.8] mag (these magnitudes are measured in circular apertures of radius equal to the PSF FWHM, corrected to infinite extent). The detection fractions depend strongly on F444W magnitude. For the brightest galaxies, F444W26.5 mag, the detection fractions are (in %) [93,100,73,47,20,13,7,0] in [F560W,F770W,F1000W,F1280W,F1500W,F1800W, F2100W,F2550W], respectively.
We divided the sample of LRDs into 4 subsets detected up to a given MIRI band: (1) 5 galaxies detected in F1280W and longer wavelengths, hereafter the Golden Five; (2) 7 galaxies detected only up to F1000W (i.e., excluding the previous type); (3) 7 sources detected only up to F770W (excluding the previous types); (4) and, lastly, the 12 galaxies not detected in any MIRI band. The panels and histograms in Figure 1, as well as the postage stamps in Figure 2, show these subsets with different colors. The first figure illustrates that the MIRI detection fraction depends strongly on the F444W magnitude and it is nearly independent of the F150W-F200W color. Interestingly, the F277W-F444W color does exhibit some differences between subsets, with the Golden Five being among the bluest, and the F1000W sample among the reddest (we will comment on this in the following sections). It is worth mentioning that only two of the Golden Five galaxies would be identified as LRDs under a more restrictive color selection of F277W-F444W1.5 mag.